# Superconductivity in spin-orbit coupled BaBi3 formed by in situ reduction of bismuthate films

### v1

@@ Conclusion
In conclusion, we observe quasi-two-dimensional superconductivity at 6 K in in situ reduced perovskite BaBiO3 films. The structural characterization of the superconducting films points to the decomposition of the reduced perovskite phase to form the superconducting BaBi3 phase. Although BaBi3 has been studied in bulk crystals due to the role of spin-orbit coupling in enhancing superconductivity, realizing superconductivity in films with reduced dimensionality allows investigating the effects of confinement and proximity effects with other bismuth-based layered topological quantum materials[53, 54].
II Materials and Methods
@@
Thin films of perovskite-phase BaBiO3 were grown to thickness of 60-90 unit cells (uc) on 5 mm × 5 mm SrTiO3 (001) substrates (commercially available) using molecular beam epitaxy (MBE). Before thin film synthesis, the substrates were cleaned at 575oC in an activated radio frequency oxygen plasma at a chamber pressure of 6x10-6 Torr for 15 minutes. Bi, Ba, Eu and Al were evaporated from effusion cells, with their growth rates calibrated using a quartz crystal monitor. The p-BBO thin films were grown under the same conditions as the plasma cleaning process. After growth, the films were cooled in the oxygen plasma to 500∘C. Subsequently, the oxygen source was turned off, and 20-80 ML Eu metal were deposited in vacuum, followed by an additional 5-10 ML of Eu in an oxygen partial pressure of 2x10-9 Torr, as measured by a residual gas analyzer. Next, 10 ML of EuO was deposited in an oxygen partial pressure of 1x10-9 Torr. The films were then cooled to room temperature for the in situ deposition of Al, forming an AlOx capping layer. Finally, the samples were transferred for ex-situ X-ray diffraction and transport measurements.
@@
Cross-sectional STEM samples of the films were prepared by using a Ga+ ion beam in a Thermo Scientific Scios 2 DualBeam system. A probe aberration corrected Thermo Fisher Titan 80-300 kV microscope operating at 200 kV was used for STEM HAADF imaging and EDX measurements.
Acknowledgements.
D.P.K. and S. acknowledge support from the Duke Endowment Duke Science and Technology Fund. Transport measurements by J.T.M. and G.F. were supported by the National Science Foundation grant DMR-2327535. STEM imaging was performed in part at the Analytical Instrumentation Facility (AIF) at North Carolina State University, which is supported by the State of North Carolina and the National Science Foundation (award number ECCS-2025064). The AIF is a member of the North Carolina Research Triangle Nanotechnology Network (RTNN), a site in the National Nanotechnology Coordinated Infrastructure (NNCI).References
- [1] A Wl Sleight, JL Gillson, and PE Bierstedt. High-temperature superconductivity in the BaPb1-xBixO3 systems. Solid State Communications, 17(1):27–28, 1975.
- [2] LF Mattheiss, EM Gyorgy, and DW Johnson Jr. Superconductivity above 20 K in the Ba-K-Bi-O system. Physical Review B, 37(7):3745, 1988.
- [3] RJ Cava, Ba Batlogg, JJ Krajewski, Re Farrow, LW Rupp Jr, AE White, K Short, WF Peck, and T Kometani. Superconductivity near 30 k without copper: the Ba0.6K0.4BiO3 perovskite. Nature, 332(6167):814–816, 1988.
- [4] DG Hinks, B Dabrowski, JD Jorgensen, AW Mitchell, DR Richards, Shiyou Pei, and Donglu Shi. Synthesis, structure and superconductivity in the Ba1-x KxBiO3-y system. Nature, 333(6176):836–838, 1988.
- [5] Shiyou Pei, JD Jorgensen, B Dabrowski, DG Hinks, DR Richards, AW Mitchell, JM Newsam, SK Sinha, D Vaknin, and AJ Jacobson. Structural phase diagram of the Ba1-xKxBiO3 system. Physical Review B, 41(7):4126, 1990.
- [6] DE Cox and AW Sleight. Crystal structure of Ba2Bi3+Bi5+O6. Solid State Communications, 19(10):969–973, 1976.
- [7] Arthur W Sleight. Bismuthates: BaBiO3 and related superconducting phases. Physica C: Superconductivity and its Applications, 514:152–165, 2015.
- [8] Kateryna Foyevtsova, Arash Khazraie, Ilya Elfimov, and George A Sawatzky. Hybridization effects and bond disproportionation in the bismuth perovskites. Physical Review B, 91(12):121114, 2015.
- [9] DF Shao, X Luo, WJ Lu, L Hu, XD Zhu, WH Song, XB Zhu, and YP Sun. Spin-orbit coupling enhanced superconductivity in bi-rich compounds (A= Sr and Ba). Scientific Reports, 6(1):21484, 2016.
- [10] Lewis Powell, Wenjun Kuang, Gabriel Hawkins-Pottier, Rashid Jalil, John Birkbeck, Ziyi Jiang, Minsoo Kim, Yichao Zou, Sofiia Komrakova, Sarah Haigh, et al. Multiphase superconductivity in PdBi2. Nature Communications, 16(1):291, 2025.
- [11] BT Matthias and JK Hulm. A search for new superconducting compounds. Physical Review, 87(5):799, 1952.
- [12] Neel Haldolaarachchige, SK Kushwaha, Quinn Gibson, and RJ Cava. Superconducting properties of BaBi3. Superconductor Science and Technology, 27(10):105001, 2014.
- [13] Rajveer Jha, Marcos A Avila, and Raquel A Ribeiro. Hydrostatic pressure effect on the superconducting properties of BaBi3 and SrBi3 single crystals. Superconductor Science and Technology, 30(2):025015, 2016.
- [14] Yanan Wang, Tomoya Taguchi, Huan Li, Ai Suzuki, Yanting Zhang, Akari Miura, Mitsuki Ikeda, Hidenori Goto, Ritsuko Eguchi, Takafumi Miyazaki, et al. Superconducting properties of BaBi3 at ambient and high pressures. Physical Chemistry Chemical Physics, 23(40):23014–23023, 2021.
- [15] N Haberkorn, Raquel A Ribeiro, Li Xiang, SL Bud’ko, and Paul C Canfield. Superconducting properties and vortex pinning in intermetallic BaBi3 single crystals: a magnetization study. Physica C: Superconductivity and its Applications, 615:1354387, 2023.
- [16] Wen-Lin Wang, Yi-Min Zhang, Nan-Nan Luo, Jia-Qi Fan, Chong Liu, Zi-Yuan Dou, Lili Wang, Wei Li, Ke He, Can-Li Song, et al. Surface symmetry breaking and disorder effects on superconductivity in perovskite BaBi3 epitaxial films. Physical Review B, 98(6):064511, 2018.
- [17] Haoxiong Zhang, Awabaikeli Rousuli, Kenan Zhang, Laipeng Luo, Chenguang Guo, Xin Cong, Zuzhang Lin, Changhua Bao, Hongyun Zhang, Shengnan Xu, et al. Tailored ising superconductivity in intercalated bulk NbSe2. Nature Physics, 18(12):1425–1430, 2022.
- [18] Gang Li, Binghai Yan, Ronny Thomale, and Werner Hanke. Topological nature and the multiple dirac cones hidden in bismuth high-Tc superconductors. Scientific Reports, 5(1):10435, 2015.
- [19] Binghai Yan, Martin Jansen, and Claudia Felser. A large-energy-gap oxide topological insulator based on the superconductor BaBiO3. Nature Physics, 9(11):709–711, 2013.
- [20] Weiliang Qiao, Jiali Zhao, Yuxiao Chen, Shenggen Cao, Wenyu Xing, Ranran Cai, Liangliang Guo, Tian Qian, XC Xie, and Wei Han. Observation of quasi-two-dimensional superconductivity at the EuO-BaBiO3 interface. Physical Review B, 109(5):054509, 2024.
- [21] Wei Guo, Agham B Posadas, Sirong Lu, David J Smith, and Alexander A Demkov. Euo epitaxy by oxygen scavenging on SrTiO3 (001): Effect of SrTiO3 thickness and temperature. Journal of Applied Physics, 124(23):235301, 2018.
- [22] Thomas Mairoser, Julia A Mundy, Alexander Melville, Daniel Hodash, Paul Cueva, Rainer Held, Artur Glavic, Jürgen Schubert, David A Muller, Darrell G Schlom, et al. High-quality EuO thin films the easy way via topotactic transformation. Nature Communications, 6(1):7716, 2015.
- [23] Robert J Zollweg. X-ray lattice constant of barium oxide. Physical Review, 100(2):671, 1955.
- [24] BT Matthias, RM Bozorth, and JH Van Vleck. Ferromagnetic interaction in EuO. Physical Review Letters, 7(5):160, 1961.
- [25] Jie Yuan, Qihong Chen, Kun Jiang, Zhongpei Feng, Zefeng Lin, Heshan Yu, Ge He, Jinsong Zhang, Xingyu Jiang, Xu Zhang, Yujun Shi, Yanmin Zhang, Mingyang Qin, Zhi Gang Cheng, Nobumichi Tamura, Yi-feng Yang, Tao Xiang, Jiangping Hu, Ichiro Takeuchi, Kui Jin, and Zhongxian Zhao. Scaling of the strange-metal scattering in unconventional superconductors. Nature, 602:431-436, 2022.
- [26] Alekseĭ Alekseevich Abrikosov. Fundamentals of the Theory of Metals. NorthHolland, Amsterdam, 1988.
- [27] HM Jaeger, DB Haviland, BG Orr, and AM Goldman. Onset of superconductivity in ultrathin granular metal films. Physical Review B, 40(1):182, 1989.
- [28] Patrick Lömker and Martina Müller. Redox-controlled epitaxy and magnetism of oxide heterointerfaces: EuO/SrTiO3. Physical Review Materials, 3(6):061401, 2019.
- [29] Kristy J Kormondy, Lingyuan Gao, Xiang Li, Sirong Lu, Agham B Posadas, Shida Shen, Maxim Tsoi, Martha R McCartney, David J Smith, Jianshi Zhou, et al. Large positive linear magnetoresistance in the two-dimensional t2g electron gas at the EuO/SrTiO3 interface. Scientific Reports, 8(1):7721, 2018.
- [30] K Zou, Sohrab Ismail-Beigi, Kim Kisslinger, Xuan Shen, Dong Su, FJ Walker, and CH Ahn. LaTiO3/KTaO3 interfaces: A new two-dimensional electron gas system. APL Materials, 3(3), 2015.
- [31] Hongrui Zhang, Yu Yun, Xuejing Zhang, Hui Zhang, Yang Ma, Xi Yan, Fei Wang, Gang Li, Rui Li, Tahira Khan, et al. High-mobility spin-polarized two-dimensional electron gases at EuO/KTaO3 interfaces. Physical Review Letters, 121(11):116803, 2018.
- [32] Nand Kumar, Neha Wadehra, Ruchi Tomar, Shama, Sanjeev Kumar, Yogesh Singh, Sushanta Dattagupta, and Suvankar Chakraverty. Observation of shubnikov-de haas oscillations, planar hall effect, and anisotropic magnetoresistance at the conducting interface of EuO–KTaO3. Advanced Quantum Technologies, 4(1):2000081, 2021.
- [33] Changjiang Liu, Xi Yan, Dafei Jin, Yang Ma, Haw-Wen Hsiao, Yulin Lin, Terence M Bretz-Sullivan, Xianjing Zhou, John Pearson, Brandon Fisher, et al. Two-dimensional superconductivity and anisotropic transport at KTaO3 (111) interfaces. Science, 371(6530):716–721, 2021.
- [34] M Tinkham. Introduction to superconductivity, Mac-Graw-Hill Inc. New York, London, Tokyo, 1996.
- [35] JM Lu, O Zheliuk, Inge Leermakers, Noah FQ Yuan, Uli Zeitler, Kam Tuen Law, and JT Ye. Evidence for two-dimensional ising superconductivity in gated MoS2. Science, 350(6266):1353–1357, 2015.
- [36] Wenzheng Wei, Dung Vu, Zhan Zhang, Frederick J Walker, and Charles H Ahn. Superconducting Nd1-xEuxNiO2 thin films using in situ synthesis. Science Advances, 9(27):eadh3327, 2023.
- [37] Athby H Al-Tawhid, Samuel J Poage, Salva Salmani-Rezaie, Antonio Gonzalez, Shalinee Chikara, David A Muller, Divine P Kumah, Maria N Gastiasoro, José Lorenzana, and Kaveh Ahadi. Enhanced critical field of superconductivity at an oxide interface. Nano Letters, 23(15):6944–6950, 2023.
- [38] Wenjie Sun, Yueying Li, Ruxin Liu, Jiangfeng Yang, Jiayi Li, Wei Wei, Gangjian Jin, Shengjun Yan, Haoying Sun, Wei Guo, et al. Evidence for anisotropic superconductivity beyond pauli limit in infinite-layer lanthanum nickelates. Advanced Materials, 35(32):2303400, 2023.
- [39] BS Chandrasekhar. A note on the maximum critical field of high-field superconductors. Appl. Phys. Letters, 1:7–8, 1962.
- [40] Albert M Clogston. Upper limit for the critical field in hard superconductors. Physical Review Letters, 9(6):266, 1962.
- [41] Yu Saito, Yasuharu Nakamura, Mohammad Saeed Bahramy, Yoshimitsu Kohama, Jianting Ye, Yuichi Kasahara, Yuji Nakagawa, Masaru Onga, Masashi Tokunaga, Tsutomu Nojima, et al. Superconductivity protected by spin–valley locking in ion-gated MoS2. Nature Physics, 12(2):144–149, 2016.
- [42] MR Beasley, JE Mooij, and TP Orlando. Possibility of vortex-antivortex pair dissociation in two-dimensional superconductors. Physical Review Letters, 42(17):1165, 1979.
- [43] K Epstein, AM Goldman, and AM Kadin. Vortex-antivortex pair dissociation in two-dimensional superconductors. Physical Review Letters, 47(7):534, 1981.
- [44] B. I. Halperin and David R. Nelson. Resistive transition in superconducting films. J Low Temp Phys, 36:599-616, 1979.
- [45] Hui Cao, Hongli Guo, Yu-Cheng Shao, Qixin Liu, Xuefei Feng, Qinwen Lu, Zhongping Wang, Aidi Zhao, Atsushi Fujimori, Yi-De Chuang, et al. Realization of electron antidoping by modulating the breathing distortion in BaBiO3. Nano Letters, 21(9):3981–3988, 2021.
- [46] P Lightfoot, Joe A Hriljac, Shiyou Pei, Ying Zheng, AW Mitchell, DR Richards, B Dabrowski, JD Jorgensen, and DG Hinks. BaBiO2.5, a new bismuth oxide with a layered structure. Journal of Solid State Chemistry, 92(2):473–479, 1991.
- [47] Danfeng Li, Kyuho Lee, Bai Yang Wang, Motoki Osada, Samuel Crossley, Hye Ryoung Lee, Yi Cui, Yasuyuki Hikita, and Harold Y Hwang. Superconductivity in an infinite-layer nickelate. Nature, 572(7771):624–627, 2019.
- [48] Matthias Hepting, Danfeng Li, CJ Jia, Haiyu Lu, E Paris, Y Tseng, X Feng, M Osada, E Been, Y Hikita, et al. Electronic structure of the parent compound of superconducting infinite-layer nickelates. Nature Materials, 19(4):381–385, 2020.
- [49] Shengwei Zeng, Changjian Li, Lin Er Chow, Yu Cao, Zhaoting Zhang, Chi Sin Tang, Xinmao Yin, Zhi Shiuh Lim, Junxiong Hu, Ping Yang, et al. Superconductivity in infinite-layer nickelate La1-xCaxNiO2 thin films. Science Advances, 8(7):eabl9927, 2022.
- [50] Yusuke Nomura and Ryotaro Arita. Superconductivity in infinite-layer nickelates. Reports on Progress in Physics, 85(5):052501, 2022.
- [51] Yaning Cui, Heng Gao, Yongchang Li, Shaowen Xu, Haotian Wang, and Wei Ren. Spin-orbit coupling enhanced electron-phonon superconductivity in infinite-layer BaBiO2. AIP Advances, 13(12), 2023.
- [52] Oleksandr I Malyi, Julien Varignon, and Alex Zunger. Bulk ndnio2 is thermodynamically unstable with respect to decomposition while hydrogenation reduces the instability and transforms it from metal to insulator. Physical Review B, 105(1):014106, 2022.
- [53] Yuqi Xia, Dong Qian, David Hsieh, L Wray, Arijeet Pal, Hsin Lin, Arun Bansil, DHYS Grauer, Yew San Hor, Robert Joseph Cava, et al. Observation of a large-gap topological-insulator class with a single dirac cone on the surface. Nature Physics, 5(6):398–402, 2009.
- [54] Shuichi Murakami. Quantum spin hall effect and enhanced magnetic response by spin-orbit coupling. Physical Review Letters, 97(23):236805, 2006.

### v2

@@ Conclusion
In conclusion, we observe superconductivity at 6 K in in situ reduced perovskite BaBiO3 films. The structural characterization of the superconducting films points to the decomposition of the reduced perovskite phase to form the superconducting BaBi3 phase. Although BaBi3 has been studied in bulk crystals due to the role of spin-orbit coupling in enhancing superconductivity, realizing superconductivity in films with reduced dimensionality allows investigating the effects of confinement and proximity effects with other bismuth-based layered topological quantum materials[61, 62].
Acknowledgments
D.P.K. and S. acknowledge support from the NSF DMR 2324174 and the Duke Endowment Duke Science and Technology Fund. Transport measurements by J.T.M. and G.F. were supported by the National Science Foundation grant DMR-2327535. STEM imaging was performed at the Analytical Instrumentation Facility (AIF) at North Carolina State University, which is supported by the State of North Carolina and the National Science Foundation (award number ECCS-2025064). The AIF is a member of the North Carolina Research Triangle Nanotechnology Network (RTNN), a site in the National Nanotechnology Coordinated Infrastructure (NNCI).
Competing interests:
There are no competing interests to declare.
Data and materials availability:
All data are available in the main text or the supplementary materials.
Acknowledgments
D.P.K. and S. acknowledge support from the NSF (Grant No. DMR 2324174) and the Duke Endowment Duke Science and Technology Fund. Transport measurements by J.T.M. and G.F. were supported by the National Science Foundation grant DMR-2327535. STEM imaging was performed at the Analytical Instrumentation Facility (AIF) at North Carolina State University, which is supported by the State of North Carolina and the National Science Foundation (award number ECCS-2025064). The AIF is a member of the North Carolina Research Triangle Nanotechnology Network (RTNN), a site in the National Nanotechnology Coordinated Infrastructure (NNCI).
Author contributions:
Conceptualization: DPK Methodology: DPK, GF Investigation: S, DPK, GJ, JTM, MB Visualization: S, DPK, JTM Supervision: DPK, GF Writing—original draft: DPK Writing—review and editing: S, DPK, GF, JTM
Competing interests:
There are no competing interests to declare.
Data and materials availability:
All data are available in the main text or the supplementary materials.
Supplementary materials
Materials and Methods
Supplementary text
Figs. S1 to S6
Supplementary Materials
Shama1,
Jordan T. McCourt1,
Merve Baksi1,Gleb Finkelstein1, &
Divine Kumah1∗
1Department of Physics, Duke University, Durham, NC, 27708, U.S.A.
∗Corresponding author. Email: divine.kumah@duke.edu
This PDF file includes:
Materials and Methods
Figure S1
Figure S2
Figure S3
Figure S4
Figure S5
Figure S6
Materials and Methods
@@ Acknowledgments
Thin films of perovskite-phase BaBiO3 were grown to thickness of 60-90 unit cells (uc) on 5 mm × 5 mm SrTiO3 (001) substrates (commercially available) using molecular beam epitaxy (MBE). Before thin film synthesis, the substrates were cleaned at 575oC in an activated radio frequency oxygen plasma at a chamber pressure of 6x10-6 Torr for 15 minutes. Bi, Ba, Eu and Al were evaporated from effusion cells, with their growth rates calibrated using a quartz crystal monitor. The p-BBO thin films were grown under the same conditions as the plasma cleaning process. After growth, the films were cooled in the oxygen plasma to 500∘C. Subsequently, the oxygen source was turned off, and 20-80 ML Eu metal were deposited in vacuum, followed by an additional 5-10 ML of Eu in an oxygen partial pressure of 2x10-9 Torr, as measured by a residual gas analyzer. Next, 10 ML of EuO was deposited in an oxygen partial pressure of 1x10-9 Torr. The films were then cooled to room temperature for in situ deposition of Al, forming an AlOx capping layer. Finally, the samples were transferred for ex-situ X-ray diffraction and transport measurements. A second set of samples were fabricated without Eu. Al was deposited at 500 oC in vacuum and the samples were subsequently cooled to room temperature to form nominal Al/BaBiO3 bilayers.
@@
Cross-sectional STEM samples of the films were prepared using a Ga+ ion beam in a Thermo Scientific Scios 2 dual beam system. A probe aberration-corrected Thermo Fisher Titan 80-300 kV microscope operating at 200 kV was used for STEM HAADF imaging and EDX measurements.
Transport measurements were performed by conventional four-probe techniques using Al wires ultrasonically bonded to the samples. The measurements were performed in a Quantum Design Dynacool PPMS system.